# Towards Adapting Open-Source Large Language Models for Expert-Level Clinical Note Generation

### v5

@@ A.3.5 RLAIF Experiment 3
Approach: At this stage, our efforts were concentrated on a limited LR search for DistillDirect. We evaluated an LR of 5e-6 and 5e-7 and compared the effects of 1 epoch versus 3 epochs in each round across three rounds. After each round of training, we resampled outputs from the updated model checkpoint to serve as “rejected” samples for the subsequent round of DPO, while continuing to use notes produced by Gemini-pro as the “preferred” samples.
@@ Results:
    Appendix Figure 7: Example Training Set Accuracy and Reward Margin during DPO with a LR of 5e-7. Examples taken from 13B_chat_short_R1. Training set accuracy reaches 0.98 at the end of 2nd epoch. After 3 epochs, the reward margin only reaches a level of 0.25.
    Appendix Figure 8: Example Training Set Accuracy and Reward Margin During DPO with a LR of 5e-6. Examples taken from 13B_chat_short_R1. Training set accuracy reaches 0.98 at the end of 2nd epoch. After 3 epochs, the reward margin only reaches a level of 0.25.
@@
  • Impact of Temperature: Consistent with previous experiments, model performance improves when a lower temperature was applied during generation (see Appendix Table 6). This observation motivated us to use lower temperature for RLHF generations.
Solution: We identified the LR of 5e-6 with a 1-epoch per round as the stable setup for DistillDirect in our experiment. Upon manual examination of outputs from all model checkpoints within this configuration, we selected the 13B_chat_short model from round 3 for the next stage of RLHF (Appendix Table 5).
@@
 13B_short 0.2463 0.5124 0.4820 0.5310 0.2277 0.5123 0.4547 0.4808
 13B_long 0.2525 0.4406 0.5423 0.2974 0.2565 0.4652 0.4984 0.3316
 13B_chat_short 0.3475 0.4633 0.4727 0.5422 0.3055 0.5293 0.5012 0.5086
 13B_chat_long 0.3463 0.4936 0.5097 0.5327 0.3620 0.5290 0.5008 0.5247
@@
 13B_short 0.2463 0.3708 0.4794 0.4775 0.2277 0.4318 0.4995 0.4956
 13B_long 0.2525 0.3730 0.4427 0.4494 0.2565 0.3609 0.4750 0.4578
 13B_chat_short 0.3475 0.4254 0.4804 0.4878 0.3055 0.4329 0.5140 0.5182
 13B_chat_long 0.3463 0.4135 0.4600 0.4601 0.3620 0.4642 0.4843 0.4662
@@
 13B_short 0.2463 0.2546 0.2655 0.2744 0.2277 0.2618 0.2573 0.3002
 13B_long 0.2525 0.2151 0.2382 0.2115 0.2565 0.2458 0.3149 0.2755
 13B_chat_short 0.3475 0.3518 0.3718 0.3428 0.3055 0.3398 0.3449 0.3840
 13B_chat_long 0.3463 0.3311 0.2984 0.3296 0.3620 0.3310 0.3719 0.3747
@@
 13B_short 0.2463 0.2864 0.2921 0.2644 0.2277 0.2599 0.2731 0.2735
 13B_long 0.2525 0.2600 0.2333 0.1839 0.2565 0.2473 0.2216 0.2453
 13B_chat_short 0.3475 0.3117 0.3388 0.3356 0.3055 0.3303 0.3370 0.3231
 13B_chat_long 0.3463 0.3090 0.2943 0.3134 0.3620 0.3174 0.3194 0.3693
@@ A.4 RLHF
 13B_short 0.4006 0.5165 0.5143 0.5349 0.4032 0.5471 0.4996 0.5246
 13B_long 0.3570 0.5049 0.5679 0.4289 0.4298 0.5064 0.5102 0.3976
 13B_chat_short 0.4390 0.5343 0.5309 0.5614 0.4825 0.5524 0.5266 0.5575
 13B_chat_long 0.4272 0.4815 0.5759 0.5605 0.4459 0.5299 0.5510 0.5294
@@
 13B_short 0.4006 0.4971 0.5145 0.5200 0.4032 0.4950 0.5464 0.5222
 13B_long 0.3570 0.4449 0.5126 0.5000 0.4298 0.5123 0.5190 0.5285
 13B_chat_short 0.4390 0.5155 0.5167 0.5352 0.4825 0.5380 0.5392 0.5411
 13B_chat_long 0.4272 0.4863 0.4980 0.5077 0.4459 0.5094 0.5255 0.5346
@@
 13B_short 0.4006 0.3760 0.3771 0.4573 0.4032 0.4390 0.4281 0.4315
 13B_long 0.3570 0.4022 0.3909 0.3693 0.4298 0.4154 0.4428 0.4411
 13B_chat_short 0.4390 0.4434 0.4482 0.4394 0.4825 0.4761 0.4625 0.4833
 13B_chat_long 0.4272 0.4398 0.4058 0.4282 0.4459 0.4763 0.4633 0.4545
@@
 13B_short 0.4006 0.4241 0.4112 0.4140 0.4032 0.4345 0.4172 0.4107
 13B_long 0.3570 0.4099 0.4175 0.4039 0.4298 0.3671 0.3865 0.4285
 13B_chat_short 0.4390 0.4291 0.4541 0.4337 0.4825 0.4569 0.4826 0.4450
 13B_chat_long 0.4272 0.4188 0.4048 0.3986 0.4459 0.4606 0.4552 0.4446
@@ B.1 ACI-BENCH Subsets
ACI-BENCH comprises five data subsets: train, validation, test1, test2 and test3 yim2023aci. Test1 and test2 corresponds to the test sets from ACL ClinicalNLP MEDIQA-Chat 2023 TaskB and TaskC, respectively abacha2023overview. Test3 corresponds to TaskC of CLEF MEDIQA-SUM 2023 yim2023overview. Given the scarcity of publicly available clinical dialogue-note datasets, we used the train, test1, and test3 subsets for various stages of model development in our study. The blinded clinical reader study was performed on the subset of test1.
@@ B.2 MIMIC-IV
The publicly available MIMIC-IV dataset comprises 431,231 unique hospital admissions from 299,712 patients admitted to an ICU or the ED of the Beth Israel Deaconess Medical Center in Boston, Massachusetts johnson2023mimic. MIMIC-IV is deidentified according to the Health Insurance Portability and Accountability Act (HIPAA) Safe Harbor provision johnson2023mimic. Access to MIMIC-IV can be requested at https://physionet.org/content/mimiciv/, which requires a signed safe usage agreement.
@@ C.1 LoRA
We used LoRA to train LLaMA-2-13B models for all phases of training. LoRA is a method that involves freezing the pre-trained model weights and only training a small percentage (<1%) of the model weights, i.e., by incorporating trainable rank decomposition matrices into each layer of the transformer architecture hu2021lora. As a quick summary, let us assume that we have the original weight matrix $\bm{W_{0}}\in\mathbb{R}^{d\times k}$. LoRA works by adding a low-rank matrix to the original weight matrix: $\Delta\bm{W}+\bm{W_{0}},\Delta\bm{W}=\bm{B}\bm{A}$ where $\bm{B}\in\mathbb{R}^{d\times r}$ and $\bm{A}\in\mathbb{R}^{r\times k}$. $r<<d$, so the matrices $\bm{B},\bm{A}$ are limited by a lower rank $r$, reducing the need to train all the parameters. Training is only performed on this $\Delta\bm{W}$, and original model weights are kept the same. We then scale $\Delta\bm{W}$ by $\frac{\alpha}{r}$, where $\alpha$ is a constant in $r$.
@@
C.2 Continued Pretraining
We followed the training scripts outlined in Meta’s official LLaMA recipe repository LlamaRecipes. We employed mixed-precision training with a batching strategy of packing and a context length of 4096 tokens. We utilized Fully Sharded Data Parallel (FSDP) on either 4 Nvidia A6000 or 4 Nvidia A100 GPUs. We maintained a batch size of 4 during training with a gradient accumulation step of 1. Consistent with LLaMA-2, we set a peak learning rate of 3e-4 for the continued pre-training stage. The AdamW optimizer with a cosine learning rate scheduler was used, and the model was trained for one epoch. The exponential moving average of training loss as shown in Figure 3 was calculated using the pracma package from R with a window size of 250.
C.3 SFT
We used a similar experiment setup as continued pretraining, including following LLaMA-recipes to perform mixed precision training on 4 GPUs using FSDP. For SFT, we selected the batching strategy of padding and trained on 3 epochs. Consistent with LLaMA-2, we set a peak learning rate of 2e-5. We truncate prompt (including instruction and dialogue) at a max length of 3000 tokens, and truncate note to 1000 tokens. We set a value of -100 for labels on prompt tokens to zero out losses from prompts.
C.4 RLAIF
We utilized the trl library from Huggingface to conduct DistillDirect TRL2020. Due to computational limitations, experiments were conducted on a single Nvidia A100 GPU with 80GB of graphics memory. To optimize memory usage, pure BF16 training was utilized with a micro-batch size of 1 and gradient accumulation steps of 8. Following a limited learning rate search detailed in Appendix A.3 and Appendix Table 5, a learning rate of 5e-6 was chosen. The optimizer used was paged_adamw_32bit. Within the DPOTrainer class, we set the beta hyperparameter to 0.1 and passed None to ref_model. Three rounds of DistillDirect were performed, with each round involving one epoch of training.
Text generation was implemented using the Transformers library wolf2020huggingfaces. When generating “rejected” samples from the latest model checkpoint, we consistently applied do_sample=True, top_p=1.0, temperature=1.0, top_k=50, repetition_penalty=1.2, and use_fast_kernels=False. The maximum number of newly generated tokens was set to 1000.
C.5 RLHF
@@
C.6 Physician Reader Study
AI-generated notes were produced by LLaMA-Clinic and Gemini Pro, employing identical generation-related hyperparameters (temperature of 0.2, top_p of 0.7 and top_k of 40). To ensure consistency in presentation across all notes, we implemented a basic post-processing step. This step standardized aspects like style, spacing, and capitalization to minimize any formatting discrepancies between human-authored and model-generated notes.
Three licensed physician evaluators, who specialize in general internal medicine or family medicine, boast rich experience in outpatient practice. The notes were presented in a random order, anonymized to remove any identifying information, and labeled as note 1, note 2, and note 3 to mask the origin of each note from the evaluators. Before assessing the notes, evaluators were instructed to read the entire patient-provider conversation. They were then asked to rate the quality of each note across three criteria: “accuracy,” “completeness,” and “real-world readiness.” For each criterion, a scoring system from 1 to 5 was used, ranging from very poor to very good, with higher scores reflecting superior quality. Specifically for “real-world readiness,” evaluators were prompted to consider the scenario of integrating AI-generated clinical notes into their daily practice, including the necessity to proofread and potentially edit these notes before filing them into medical records.
C.7 Statistical Analysis
The non-parametric Kruskal-Wallis H test was selected to compare differences in word counts among the three-note groups, utilizing the scipy package in Python virtanen2020scipy. We measured IRR using Gwet’s AC2 statistics implemented through the irrCAC package in R gwet2014handbook; gwet2019package. We reported results with quadratic weights for Gwet’s AC2, as this approach is reliable for ordinal data against the grey zones tran2021impact. Due to our relatively small reviewer pool, we chose not to conduct statistical significance testing on the physician reader study, aligning with practices observed in the deep learning community touvron2023llama; stiennon2020learning.
C.8 Model Development Cost Estimation
@@
C.9 Model Inference Cost Estimation
@@
For open-source models, we assumed deployment on Fireworks.ai, a company that offers serverless inference for customized LLMs. We sourced pricing information from the websites of Google AI, OpenAI, and Fireworks.ai in May 2024 for on-demand API calls. Detailed pricing information can be found in Appendix Table 12. We assumed an average of 3,000 input tokens and 1,000 output tokens per request for clinical note generation. This estimation likely contains redundancy and leaves room for prompt engineering, given that in a cohort of real-world family medicine clinical encounters, the average lengths per dialogue and note are 1505 and 683 tokens, respectively yim2023aci.
An important consideration for production is ensuring adequate throughput for LLMs. As an example, Gemini 1.0 Pro has a rate limit of 360 requests per minute, while LLaMA-Clinic, deployed with the "Developer" plan from Fireworks.ai, allows 600 requests per minute. We consider this default rate limit acceptable for our calculations, as shown in Figure 5. For example, 1 million annual requests translate into an average of approximately 5.7 requests per minute (RPM), using the formula:
@@
assuming an 8-hour workday. However, this calculation does not account for peak demand, which would necessitate system redundancy in a production environment. In addition, there are other technical factors to consider for deployment in production, such as latency and throughput variance ArtificialAnalysisAI2024, which were not included in our analysis.

### v6

@@ A.3.5 RLAIF Experiment 3
Approach: We evaluated an LR of 5e-6 and 5e-7 and compared the effects of 1 epoch versus 3 epochs in each round across three rounds. After each round of training, we resampled outputs from the updated model checkpoint to serve as “rejected” samples for the subsequent round of DPO, while continuing to use notes produced by Gemini-pro as the “preferred” samples.
@@ Results:
    Appendix Figure 3: Example Training Set Accuracy and Reward Margin during DPO with a LR of 5e-7. Examples taken from 13B-chat-short_R1. Training set accuracy reaches 0.98 at the end of 2nd epoch. After 3 epochs, the reward margin only reaches a level of 0.25.
    Appendix Figure 4: Example Training Set Accuracy and Reward Margin During DPO with a LR of 5e-6. Examples taken from 13B-chat-short_R1. Training set accuracy reaches 0.98 at the end of 2nd epoch. After 3 epochs, the reward margin only reaches a level of 0.25.
@@
  • Impact of Temperature: Consistent with previous experiments, model performance improves when a lower temperature was applied during generation (see Appendix Table 5).
Solution: We identified the LR of 5e-6 with a 1-epoch per round as the stable setup for DistillDirect in our experiment. Upon manual examination of outputs from all model checkpoints within this configuration, we selected the 13B-chat-short model from round 3 for the next stage of RLHF (Appendix Table 4).
@@
 13B-short 0.2463 0.5124 0.4820 0.5310 0.2277 0.5123 0.4547 0.4808
 13B-long 0.2525 0.4406 0.5423 0.2974 0.2565 0.4652 0.4984 0.3316
 13B-chat-short 0.3475 0.4633 0.4727 0.5422 0.3055 0.5293 0.5012 0.5086
 13B-chat-long 0.3463 0.4936 0.5097 0.5327 0.3620 0.5290 0.5008 0.5247
@@
 13B-short 0.2463 0.3708 0.4794 0.4775 0.2277 0.4318 0.4995 0.4956
 13B-long 0.2525 0.3730 0.4427 0.4494 0.2565 0.3609 0.4750 0.4578
 13B-chat-short 0.3475 0.4254 0.4804 0.4878 0.3055 0.4329 0.5140 0.5182
 13B-chat-long 0.3463 0.4135 0.4600 0.4601 0.3620 0.4642 0.4843 0.4662
@@
 13B-short 0.2463 0.2546 0.2655 0.2744 0.2277 0.2618 0.2573 0.3002
 13B-long 0.2525 0.2151 0.2382 0.2115 0.2565 0.2458 0.3149 0.2755
 13B-chat-short 0.3475 0.3518 0.3718 0.3428 0.3055 0.3398 0.3449 0.3840
 13B-chat-long 0.3463 0.3311 0.2984 0.3296 0.3620 0.3310 0.3719 0.3747
@@
 13B-short 0.2463 0.2864 0.2921 0.2644 0.2277 0.2599 0.2731 0.2735
 13B-long 0.2525 0.2600 0.2333 0.1839 0.2565 0.2473 0.2216 0.2453
 13B-chat-short 0.3475 0.3117 0.3388 0.3356 0.3055 0.3303 0.3370 0.3231
 13B-chat-long 0.3463 0.3090 0.2943 0.3134 0.3620 0.3174 0.3194 0.3693
@@ A.4 RLHF
 13B-short 0.4006 0.5165 0.5143 0.5349 0.4032 0.5471 0.4996 0.5246
 13B-long 0.3570 0.5049 0.5679 0.4289 0.4298 0.5064 0.5102 0.3976
 13B-chat-short 0.4390 0.5343 0.5309 0.5614 0.4825 0.5524 0.5266 0.5575
 13B-chat-long 0.4272 0.4815 0.5759 0.5605 0.4459 0.5299 0.5510 0.5294
@@
 13B-short 0.4006 0.4971 0.5145 0.5200 0.4032 0.4950 0.5464 0.5222
 13B-long 0.3570 0.4449 0.5126 0.5000 0.4298 0.5123 0.5190 0.5285
 13B-chat-short 0.4390 0.5155 0.5167 0.5352 0.4825 0.5380 0.5392 0.5411
 13B-chat-long 0.4272 0.4863 0.4980 0.5077 0.4459 0.5094 0.5255 0.5346
@@
 13B-short 0.4006 0.3760 0.3771 0.4573 0.4032 0.4390 0.4281 0.4315
 13B-long 0.3570 0.4022 0.3909 0.3693 0.4298 0.4154 0.4428 0.4411
 13B-chat-short 0.4390 0.4434 0.4482 0.4394 0.4825 0.4761 0.4625 0.4833
 13B-chat-long 0.4272 0.4398 0.4058 0.4282 0.4459 0.4763 0.4633 0.4545
@@
 13B-short 0.4006 0.4241 0.4112 0.4140 0.4032 0.4345 0.4172 0.4107
 13B-long 0.3570 0.4099 0.4175 0.4039 0.4298 0.3671 0.3865 0.4285
 13B-chat-short 0.4390 0.4291 0.4541 0.4337 0.4825 0.4569 0.4826 0.4450
 13B-chat-long 0.4272 0.4188 0.4048 0.3986 0.4459 0.4606 0.4552 0.4446
@@ B.1 ACI-BENCH Subsets
ACI-BENCH comprises five data subsets: train, validation, test1, test2 and test3 Yim et al. (2023b). Test1 and test2 corresponds to the test sets from ACL ClinicalNLP MEDIQA-Chat 2023 TaskB and TaskC, respectively Abacha et al. (2023). Test3 corresponds to TaskC of CLEF MEDIQA-SUM 2023 Yim et al. (2023a). Given the scarcity of publicly available clinical dialogue-note datasets, we used the train, test2, and test3 subsets for various stages of model development in our study. The blinded clinical reader study was performed on the test1 subset.
@@ B.2 MIMIC-IV
The publicly available MIMIC-IV dataset comprises 431,231 unique hospital admissions from 299,712 patients admitted to an ICU or the ED of the Beth Israel Deaconess Medical Center in Boston, Massachusetts Johnson et al. (2023). MIMIC-IV is deidentified according to the Health Insurance Portability and Accountability Act (HIPAA) Safe Harbor provision Johnson et al. (2023). Access to MIMIC-IV can be requested at https://physionet.org/content/mimiciv/, which requires a signed safe usage agreement.
@@ C.1 LoRA
We used LoRA to train LLaMA-2-13B models for all phases of training. LoRA is a method that involves freezing the pre-trained model weights and only training a small percentage (<1%) of the model weights, i.e., by incorporating trainable rank decomposition matrices into each layer of the transformer architecture Hu et al. (2021). As a quick summary, let us assume that we have the original weight matrix $\bm{W_{0}}\in\mathbb{R}^{d\times k}$. LoRA works by adding a low-rank matrix to the original weight matrix: $\Delta\bm{W}+\bm{W_{0}},\Delta\bm{W}=\bm{B}\bm{A}$ where $\bm{B}\in\mathbb{R}^{d\times r}$ and $\bm{A}\in\mathbb{R}^{r\times k}$. $r<<d$, so the matrices $\bm{B},\bm{A}$ are limited by a lower rank $r$, reducing the need to train all the parameters. Training is only performed on this $\Delta\bm{W}$, and original model weights are kept the same. We then scale $\Delta\bm{W}$ by $\frac{\alpha}{r}$, where $\alpha$ is a constant in $r$.
@@
C.2 Medical LLMs Evaluation
We evaluated state-of-the-art medical and clinical LLMs fine-tuned on biomedical literature and clinical notes: Meditron-7B Chen et al. (2023), LLaMA-3-Med42-8B Christophe et al. (2024), and MeLLaMA-13B-chat Xie et al. (2024). The same generation configurations with one-shot prompting from our main experiments were applied, except for MeLLaMA, for which we used its default generation settings to achieve better performance.
C.3 Continued Pretraining
We followed the training scripts outlined in Meta’s official LLaMA recipe repository meta-llama . We employed mixed-precision training with a batching strategy of packing and a context length of 4096 tokens. We utilized Fully Sharded Data Parallel (FSDP) on either 4 Nvidia A6000 or 4 Nvidia A100 GPUs. We maintained a batch size of 4 during training with a gradient accumulation step of 1. Consistent with LLaMA-2, we set a peak learning rate of 3e-4 for the continued pretraining stage. The AdamW optimizer with a cosine learning rate scheduler was used, and the model was trained for one epoch. The exponential moving average of training loss as shown in Figure 3 was calculated using the pracma package from R with a window size of 250.
C.4 SFT
We used a similar experiment setup as continued pretraining, including following LLaMA-recipes to perform mixed precision training on 4 GPUs using FSDP Zhao et al. (2023). For SFT, we selected the batching strategy of padding and trained on 3 epochs. Consistent with LLaMA-2, we set a peak learning rate of 2e-5. We truncate prompt (including instruction and dialogue) at a max length of 3000 tokens, and truncate note to 1000 tokens. We set a value of -100 for labels on prompt tokens to zero out losses from prompts.
C.5 RLAIF
We utilized the trl library from Huggingface to conduct DistillDirect von Werra et al. . Due to computational limitations, experiments were conducted on a single Nvidia A100 GPU with 80GB of graphics memory. To optimize memory usage, pure BF16 training was utilized with a micro-batch size of 1 and gradient accumulation steps of 8. Following a limited learning rate search detailed in Appendix A.3 and Appendix Table 4, a learning rate of 5e-6 was chosen. The optimizer used was paged_adamw_32bit. Within the DPOTrainer class, we set the beta hyperparameter to 0.1 and passed None to ref_model. Three rounds of DistillDirect were performed, with each round involving one epoch of training.
Text generation was implemented using the Transformers library Wolf et al. (2019). When generating “rejected” samples from the latest model checkpoint, we consistently applied do_sample=True, top_p=1.0, temperature=1.0, top_k=50, repetition_penalty=1.2, and use_fast_kernels=False. The maximum number of newly generated tokens was set to 1000.
C.6 RLHF
@@
C.7 Physician Reader Study
LLM-generated notes were produced by LLaMA-Clinic and Gemini Pro, employing identical generation-related hyperparameters (temperature of 0.2, top_p of 0.7 and top_k of 40). To ensure consistency in presentation across all notes, we implemented a basic post-processing step. This step standardized aspects like style, spacing, and capitalization to minimize any formatting discrepancies between human-authored and model-generated notes.
Four licensed physician evaluators, who specialize in general internal medicine or family medicine, boast rich experience in outpatient practice. The notes were presented in a random order, anonymized to remove any identifying information, and labeled as note 1, note 2, and note 3 to mask the origin of each note from the evaluators. Before assessing the notes, evaluators were instructed to read the entire patient-provider conversation. They were then asked to rate the quality of each note across three criteria: “accuracy,” “completeness,” and “real-world readiness.” For each criterion, a scoring system from 1 to 5 was used, ranging from very poor to very good, with higher scores reflecting superior quality. Specifically for “real-world readiness,” evaluators were prompted to consider the scenario of integrating AI-generated clinical notes into their daily practice, including the necessity to proofread and potentially edit these notes before filing them into medical records.
C.8 Statistical Analysis
The non-parametric Kruskal-Wallis H test was selected to compare differences in word counts among the three-note groups, utilizing the scipy package in Python Virtanen et al. (2020). We measured IRR using Gwet’s AC2 statistics implemented through the irrCAC package in R Gwet (2014); Gwet and Gwet (2019). We reported results with quadratic weights for Gwet’s AC2, as this approach is reliable for ordinal data against the grey zones Tran et al. (2021). Due to our relatively small reviewer pool, we chose not to conduct statistical significance testing on the physician reader study, aligning with practices observed in the deep learning community Touvron et al. (2023); Stiennon et al. (2020).
C.9 Model Development Cost Estimation
@@
C.10 Model Inference Cost Estimation
@@
For open-source models, we assumed deployment on Fireworks.ai, a company that offers serverless inference for customized LLMs. We sourced pricing information from the websites of Google AI, OpenAI, and Fireworks.ai in May 2024 for on-demand API calls. Detailed pricing information can be found in Appendix Table 11. We assumed an average of 3,000 input tokens and 1,000 output tokens per request for clinical note generation. This estimation likely contains redundancy and leaves room for prompt engineering, given that in a cohort of real-world family medicine clinical encounters, the average lengths per dialogue and note are 1505 and 683 tokens, respectively Yim et al. (2023b).
An important consideration for production is ensuring adequate throughput for LLMs. As an example, Gemini 1.0 Pro has a rate limit of 360 requests per minute, while LLaMA-Clinic, deployed with the “Developer” plan from Fireworks.ai, allows 600 requests per minute. We consider this default rate limit acceptable for our calculations, as shown in Figure 5. For example, 1 million annual requests translate into an average of approximately 5.7 requests per minute (RPM), using the formula:
@@
assuming an 8-hour workday. However, this calculation does not account for peak demand, which would necessitate system redundancy in a production environment. In addition, there are other technical factors to consider for deployment in production, such as latency and throughput variance Art (2024), which were not included in our analysis.